# Canonical bases of tensor products of integrable highest weight modules arising from framed constructions

### v2

@@ 5.1. Kashiwara’s operators on $\mathbf{f}\theta_{\lambda}\mathbf{f}$
For any $i\in I$, let $\tilde{\phi}_{i}:\tilde{\mathbf{f}}\rightarrow\tilde{\mathbf{f}},\tilde{\epsilon}_{i}:\tilde{\mathbf{f}}\rightarrow\tilde{\mathbf{f}}$ be the Kashiwara’s operators associated to the algebra $\tilde{\mathbf{f}}$, see [13, Section 17.3.1]. Recall that $\tilde{\phi}_{i}$ is induced by $\phi_{i}$ the left multiplication by $\theta_{i}$, and $\tilde{\epsilon}_{i}$ is induced by the linear map ${{}_{i}r}$, see 4.3. It is clear that $\mathbf{f}\theta_{\lambda}\mathbf{f}$ is stable under $\phi_{i}$. By (2) for the case $n=1$, we know that $\mathbf{f}\theta_{\lambda}\mathbf{f}$ is also stable under ${{}_{i}r}$. Hence $\tilde{\phi}_{i}$ and $\tilde{\epsilon}_{i}$ can be restricted to $\tilde{\phi}_{i}:\mathbf{f}\theta_{\lambda}\mathbf{f}\rightarrow\mathbf{f}\theta_{\lambda}\mathbf{f},\ \tilde{\epsilon}_{i}:\mathbf{f}\theta_{\lambda}\mathbf{f}\rightarrow\mathbf{f}\theta_{\lambda}\mathbf{f}$.
Let $\mathcal{L}(\tilde{\mathbf{f}})=\{x\in{{}_{\mathcal{A}}}\tilde{\mathbf{f}}\mid(x,x)\in\mathbf{A}\}$. By Lemma 2.4, Theorem 2.5 and $\mathcal{A}\cap\mathbf{A}=\mathbb{Z}[v^{-1}]$, we know that $\mathcal{L}(\tilde{\mathbf{f}})$ is the $\mathbb{Z}[v^{-1}]$-submodule of $\tilde{\mathbf{f}}$ generated by $\tilde{\mathbf{B}}$. We have the direct sum decomposition $\mathcal{L}(\tilde{\mathbf{f}})=\bigoplus_{\tilde{\nu}\in\mathbb{N}[I]}\mathcal{L}(\tilde{\mathbf{f}})_{\tilde{\nu}}$, where $\mathcal{L}(\tilde{\mathbf{f}})_{\tilde{\nu}}$ is the $\mathbb{Z}[v^{-1}]$-submodule generated by $\tilde{\mathbf{B}}_{\tilde{\nu}}$, see [13, Section 17.3.3, 18.1.5].
@@
As results, for any $i\in I$ and $b\in\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$, there exists a unique $b_{1}\in\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ such that
@@
By [13, Proposition 17.3.5, Corollary 17.3.7], we have $\tilde{\phi}_{i}(b)=\pi_{i,n+1}(b^{\prime})\ \mathrm{mod}\,v^{-1}\mathcal{L}(\tilde{\mathbf{f}})$. By Lemma 4.2, we have $\pi_{i,n+1}(b^{\prime})\in\mathcal{L}(\mathbf{f}\theta_{\lambda}\mathbf{f})$. Since $\mathcal{L}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ is stable under $\tilde{\phi}_{i}$, we have $\tilde{\phi}_{i}(b)\in\mathcal{L}(\mathbf{f}\theta_{\lambda}\mathbf{f})$. Hence $\tilde{\phi}_{i}(b)=\pi_{i,n+1}(b^{\prime})\ \mathrm{mod}\,v^{-1}\mathcal{L}(\mathbf{f}\theta_{\lambda}\mathbf{f})$. The other statements can be proved similarly. ∎
@@
We denote $\mathcal{L}^{\prime}=\sum_{\nu\in\mathbb{N}[I]}\mathcal{L}(\mathbf{f}\theta_{\lambda}\mathbf{f})^{\nu}$, then $\mathcal{L}^{\prime}$ is a submodule of $\mathcal{L}$. For any $b\in\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$, we make an induction on $\mathrm{tr}\,|b|$ to prove that there exist $b^{\prime}\in\bigcap_{i\in I}\tilde{\mathbf{B}}_{i,0}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ and $i_{1},...,i_{n}\in I$ such that $b=\tilde{\phi}_{i_{1}}...\tilde{\phi}_{i_{n}}(b^{\prime})\ \mathrm{mod}\ v^{-1}\mathcal{L}(\mathbf{f}\theta_{\lambda}\mathbf{f})$.
@@
(a) The left multiplication by $\theta_{\lambda}$ defines a map from $\mathbf{B}$ to $\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ such that $t_{i}^{\sigma}(b_{0})=t_{i}^{\sigma}(\theta_{\lambda}b_{0})$ for any $b_{0}\in\mathbf{B}$ and $i\in I$. As a result, $b_{0}\in\mathbf{B}(\xi)$ if and only if $\theta_{\lambda}b_{0}\in\tilde{\mathbf{B}}(\xi,\lambda)$.
@@
(a) For any $b_{0}\in\mathbf{B}$, it is clear that $b_{0}\in\bigcap_{i\in I}\tilde{\mathbf{B}}_{i^{\prime},0}$. For any $j\in I$, by Theorem 2.6, we have $\theta_{j^{\prime}}^{(\langle j,\lambda\rangle)}b_{0}=\pi_{j^{\prime},\langle j,\lambda\rangle}(b_{0})\in\tilde{\mathbf{B}}$. Notice that we have $\theta_{j^{\prime}}^{(\langle j,\lambda\rangle)}b_{0}\in\bigcap_{i\in I,i\not=j}\tilde{\mathbf{B}}_{i^{\prime},0}$, by repeating above process, we obtain $\theta_{\lambda}b_{0}\in\tilde{\mathbf{B}}$. By definitions, it is clear that $\theta_{\lambda}b_{0}\in\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$. It remains to prove that $t_{i}^{\sigma}(b_{0})=t_{i}^{\sigma}(\theta_{\lambda}b_{0})$ for any $i\in I$. By [13, Section 14.3.1], it is equivalent to $t_{i}(\sigma(b_{0}))=t_{i}(\sigma(\theta_{\lambda}b_{0}))=t_{i}(\sigma(b_{0})\theta_{\lambda})$ for any $i\in I$. For any $x\in\mathbf{f}$ and $i\in I$, it is clear that $\phi_{i}(x)\theta_{\lambda}=\phi_{i}(x\theta_{\lambda})$. By the property of the linear map ${{}_{i}r}$, see 4.3, we have ${{}_{i}r}(x)\theta_{\lambda}={{}_{i}r}(x\theta_{\lambda})$. Hence the right multiplication by $\theta_{\lambda}$ commutes with the actions of $\phi_{i},\epsilon_{i}$ for any $i\in I$, and so it commutes with the Kashiwara’s operators $\tilde{\phi_{i}},\tilde{\epsilon}_{i}$ on $\mathbf{f},\mathbf{f}\theta_{\lambda}\mathbf{f}$ for any $i\in I$. By [13, Corollary 17.3.7] and Lemma 5.1, we have $t_{i}(\sigma(b_{0}))=t_{i}(\sigma(b_{0})\theta_{\lambda})$ for any $i\in I$, as desired.
(b) Firstly, we prove that $\theta_{\lambda}b_{0}\in\bigcap_{i\in I}\tilde{\mathbf{B}}_{i,0}$ for any $b_{0}\in\sigma(\mathbf{B}(\lambda))$. Suppose $\theta_{\lambda}b_{0}\in\tilde{\mathbf{B}}_{j,\geqslant 1}$ for some $j\in I$. By [13, Section 14.3.1], we have $t_{j}^{\sigma}(\sigma(\theta_{\lambda}b_{0}))=t_{j}(\theta_{\lambda}b_{0})\geqslant 1$, and so $\sigma(b_{0})\theta_{\lambda}=\sigma(b_{0})\sigma(\theta_{\lambda})=\sigma(\theta_{\lambda}b_{0})\in\tilde{\mathbf{f}}\theta_{j}\subset\sum_{i\in I}\tilde{\mathbf{f}}\theta_{i}$. By Proposition 3.4, we know that $\sigma(b_{0})\in\sum_{i\in I}\mathbf{f}\theta_{i}^{\langle i,\lambda\rangle+1}$, a contradiction to $b_{0}\in\sigma(\mathbf{B}(\lambda))$. Hence $\theta_{\lambda}\cdot$ can be restricted to be a map $\sigma(\mathbf{B}(\lambda))\rightarrow\bigcap_{i\in I}\tilde{\mathbf{B}}_{i,0}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$. Secondly, for any homogeneous $x\in\mathbf{f}$, we make an induction on $\mathrm{tr}\,|x|\in\mathbb{N}$ to prove that if $\theta_{\lambda}x=0$, then $x=0$.
$\bullet$ If $\mathrm{tr}\,\nu=0$, then it is clear.
$\bullet$ If $\mathrm{tr}\,\nu>0$, for any $i\in I$, by the property of ${{}_{i}r}$, see 4.3, we have
@@
and then $\theta_{\lambda}{{}_{i}r}(x)=0$. Notice that $\mathrm{tr}\,|{{}_{i}r}(x)|=\mathrm{tr}\,|x|-1$, by the inductive hypothesis, we have ${{}_{i}r}(x)=0$, and so $x=0$ by [13, Lemma 1.2.15], as desired. Hence the map $\theta_{\lambda}\cdot:\sigma(\mathbf{B}(\lambda))\rightarrow\bigcap_{i\in I}\tilde{\mathbf{B}}_{i,0}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ is injective. Thirdly, for any $b\in\bigcap_{i\in I}\tilde{\mathbf{B}}_{i,0}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$, we make an induction on $\mathrm{tr}\,|b|$ to prove that there exists $b_{0}\in\mathbf{B}$ such that $b=\theta_{\lambda}b_{0}$.
@@
$\bullet$ If $t_{i}^{\sigma}(b)>0$ for some $i\in I$, by Theorem 2.6 and Lemma 4.2, there exists a unique $b^{\prime}\in\tilde{\mathbf{B}}_{i,0}^{\sigma}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ such that $b^{\prime}\theta_{i}^{(t_{i}^{\sigma}(b))}=b+\sum_{b^{\prime\prime}\in\tilde{\mathbf{B}}_{i,\geqslant t_{i}^{\sigma}(b)+1}^{\sigma}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})}c_{b^{\prime\prime}}b^{\prime\prime}$. We claim that $b^{\prime}\in\bigcap_{j\in I}\tilde{\mathbf{B}}_{j,0}$. Indeed, suppose that $t_{j}(b^{\prime})>0$ for some $j\in I$, Theorem 2.6 and Lemma 4.2, there exists a unique ${}^{\prime}b\in\tilde{\mathbf{B}}_{j,0}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ such that $\theta_{j}^{t_{j}(b^{\prime})}{{}^{\prime}b}=b^{\prime}+\sum_{{}^{\prime\prime}b\in\tilde{\mathbf{B}}_{j,\geqslant t_{j}(b^{\prime})}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})}{{}^{\prime}c}_{{}^{\prime\prime}b}{{}^{\prime\prime}b}$. Then we have
$\theta_{j}^{t_{j}(b^{\prime})}{{}^{\prime}b}\theta_{i}^{(t_{i}^{\sigma}(b))}=$ $b^{\prime}\theta_{i}^{(t_{i}^{\sigma}(b))}+\sum_{{}^{\prime\prime}b\in\tilde{\mathbf{B}}_{j,\geqslant t_{j}(b^{\prime})}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})}{{}^{\prime}c}_{{}^{\prime\prime}b}{{}^{\prime\prime}b}\theta_{i}^{(t_{i}^{\sigma}(b))}$
$=$ $b+\sum_{b^{\prime\prime}\in\tilde{\mathbf{B}}_{i,\geqslant t_{i}^{\sigma}(b)+1}^{\sigma}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})}c_{b^{\prime\prime}}b^{\prime\prime}+\sum_{{}^{\prime\prime}b\in\tilde{\mathbf{B}}_{j,\geqslant t_{j}(b^{\prime})}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})}{{}^{\prime}c}_{{}^{\prime\prime}b}{{}^{\prime\prime}b}\theta_{i}^{(t_{i}^{\sigma}(b))}.$
@@
Notice that $c_{b^{\prime\prime}},{{}^{\prime}c}_{{}^{\prime\prime}b}\in\mathbb{N}[v,v^{-1}]$ and ${{}^{\prime\prime}b}\theta_{i}^{(t_{i}^{\sigma}(b))}$ is a $\mathbb{N}[v,v^{-1}]$-linear combination of $\tilde{\mathbf{B}}$. Hence $b$ appears in $\theta_{j}^{t_{j}(b^{\prime})}{{}^{\prime}b}\theta_{i}^{(t_{i}^{\sigma}(b))}$. By Corollary 2.7, we have $t_{j}(b)\geqslant t_{j}(b^{\prime})>0$, a contradiction. Notice that $\mathrm{tr}\,|b^{\prime}|-\mathrm{tr}\,|b|-t_{i}^{\sigma}(b)$, by the inductive hypothesis, there exists $b^{\prime}_{0}\in\mathbf{B}$ such that $b^{\prime}=\theta_{\lambda}b^{\prime}_{0}$. By (a), we have $t_{i}^{\sigma}(b^{\prime}_{0})=t_{i}^{\sigma}(b^{\prime})=0$. By Theorem 2.6, there exists a unique $b_{0}\in\mathbf{B}_{i,t_{i}^{\sigma}(b)}^{\sigma}$ such that $b^{\prime}_{0}\theta_{i}^{(t_{i}^{\sigma}(b))}=b_{0}+\sum_{b^{\prime\prime}_{1}\in\mathbf{B}_{i,\geqslant t_{i}^{\sigma}(b)+1}^{\sigma}}c^{\prime}_{b^{\prime\prime}_{1}}b^{\prime\prime}_{1}$. Hence we have
$b-\theta_{\lambda}b_{0}=\sum_{b^{\prime\prime}_{1}\in\mathbf{B}_{i,\geqslant t_{i}^{\sigma}(b)+1}^{\sigma}}c^{\prime}_{b^{\prime\prime}_{1}}\theta_{\lambda}b^{\prime\prime}_{1}-\sum_{b^{\prime\prime}\in\tilde{\mathbf{B}}_{i,\geqslant t_{i}^{\sigma}(b)+1}^{\sigma}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})}c_{b^{\prime\prime}}b^{\prime\prime}.$
@@
By (a), we have $\theta_{\lambda}b_{0}\in\tilde{\mathbf{B}}_{i,t_{i}^{\sigma}(b)}^{\sigma}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ and $\theta_{\lambda}b^{\prime\prime}_{1}\in\tilde{\mathbf{B}}_{i,\geqslant t_{i}^{\sigma}(b)+1}^{\sigma}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$. Hence we have $b=\theta_{\lambda}b_{0}$, as desired. It remains to prove that $b_{0}\in\sigma(\mathbf{B}(\lambda))$. Suppose that $t_{j}(b_{0})\geqslant\langle j,\lambda\rangle+1$ for some $j\in I$, since $i^{\prime}\cdot j=-\delta_{i,j}$ for any $i\in I$, by the quantum Serre relations and [13, Lemma 42.1.2(b)], we have

### v3

@@ 5. Kashiwara’s operators
The purpose of this section is to prove Lemma 4.6. Our idea comes from the proof of [13, Lemma 19.1.4] which proves the almost orthonormal property of the canonical basis $\mathbf{B}(\Lambda_{\lambda})$ of $\Lambda_{\lambda}$ for any $\lambda\in X^{+}$ by using Kashiwara’s operators.
@@ 5.1. Kashiwara’s operators on $\mathbf{f}\theta_{\lambda}\mathbf{f}$
For any $i\in I$, let $\tilde{\phi}_{i}:\tilde{\mathbf{f}}\rightarrow\tilde{\mathbf{f}},\tilde{\epsilon}_{i}:\tilde{\mathbf{f}}\rightarrow\tilde{\mathbf{f}}$ be the Kashiwara’s operators associated to the algebra $\tilde{\mathbf{f}}$, see [13, Section 17.3.1]. Recall that $\tilde{\phi}_{i}$ is induced by $\phi_{i}$ the left multiplication by $\theta_{i}$, and $\tilde{\epsilon}_{i}$ is induced by the linear map $_{i}r$}, see §4.3 for the definition of $_{i}r$}. It is clear that $\mathbf{f}\theta_{\lambda}\mathbf{f}$ is stable under $\phi_{i}$. By (2) for the case $n=1$, we know that $\mathbf{f}\theta_{\lambda}\mathbf{f}$ is also stable under $_{i}r$}. Hence $\tilde{\phi}_{i}$ and $\tilde{\epsilon}_{i}$ can be restricted to $\tilde{\phi}_{i}:\mathbf{f}\theta_{\lambda}\mathbf{f}\rightarrow\mathbf{f}\theta_{\lambda}\mathbf{f},\ \tilde{\epsilon}_{i}:\mathbf{f}\theta_{\lambda}\mathbf{f}\rightarrow\mathbf{f}\theta_{\lambda}\mathbf{f}$.
[…]
@@
As a result, for any $i\in I$ and $b\in\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$, there exists a unique $b_{1}\in\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ such that
@@
[…]
@@
We denote $\mathcal{L}^{\prime}=\sum_{\nu\in\mathbb{N}[I]}\mathcal{L}(\mathbf{f}\theta_{\lambda}\mathbf{f})^{\nu}$, then $\mathcal{L}^{\prime}$ is a submodule of $\mathcal{L}$. For any $b\in\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$, we use induction on $\mathrm{tr}\,|b|$ to prove that there exist $b^{\prime}\in\bigcap_{i\in I}\tilde{\mathbf{B}}_{i,0}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ and $i_{1},...,i_{n}\in I$ such that $b=\tilde{\phi}_{i_{1}}...\tilde{\phi}_{i_{n}}(b^{\prime})\ \mathrm{mod}\ v^{-1}\mathcal{L}(\mathbf{f}\theta_{\lambda}\mathbf{f})$.
@@
(a) The left multiplication by $\theta_{\lambda}$ defines an injective map from $\mathbf{B}$ to $\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ such that $t_{i}^{\sigma}(b_{0})=t_{i}^{\sigma}(\theta_{\lambda}b_{0})$ for any $b_{0}\in\mathbf{B}$ and $i\in I$. As a result, $b_{0}\in\mathbf{B}(\xi)$ if and only if $\theta_{\lambda}b_{0}\in\tilde{\mathbf{B}}(\xi,\lambda)$.
@@
(a) Firstly, we prove that the left multiplication by $\theta_{\lambda}$ defines an injective linear map from $\mathbf{f}$ to $\tilde{\mathbf{f}}$. For any homogeneous $x\in\mathbf{f}$, we use induction on $\mathrm{tr}\,|x|\in\mathbb{N}$ to prove that if $\theta_{\lambda}x=0$, then $x=0$.
$\bullet$ If $\mathrm{tr}\,|x|=0$, then it is clear.
$\bullet$ If $\mathrm{tr}\,|x|>0$, then for any $i\in I$, by (12), we have
@@
and then $\theta_{\lambda}{{}_{i}r}(x)=0$. Notice that $\mathrm{tr}\,|{{}_{i}r}(x)|=\mathrm{tr}\,|x|-1$, by the inductive hypothesis, we have ${{}_{i}r}(x)=0$, and so $x=0$ by [13, Lemma 1.2.15], as desired. Hence the map $\theta_{\lambda}\cdot:\mathbf{f}\rightarrow\tilde{\mathbf{f}}$ is injective.
Secondly, for any $b_{0}\in\mathbf{B}$, it is clear that $b_{0}\in\bigcap_{i\in I}\tilde{\mathbf{B}}_{i^{\prime},0}$. For any $j\in I$, by Theorem 2.6, we have $\theta_{j^{\prime}}^{(\langle j,\lambda\rangle)}b_{0}=\pi_{j^{\prime},\langle j,\lambda\rangle}(b_{0})\in\tilde{\mathbf{B}}$. Notice that we have $\theta_{j^{\prime}}^{(\langle j,\lambda\rangle)}b_{0}\in\bigcap_{i\in I,i\not=j}\tilde{\mathbf{B}}_{i^{\prime},0}$, by repeating above process, we obtain $\theta_{\lambda}b_{0}\in\tilde{\mathbf{B}}$. By definitions, it is clear that $\theta_{\lambda}b_{0}\in\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$. It remains to prove that $t_{i}^{\sigma}(b_{0})=t_{i}^{\sigma}(\theta_{\lambda}b_{0})$ for any $i\in I$. By [13, Section 14.3.1], it is equivalent to $t_{i}(\sigma(b_{0}))=t_{i}(\sigma(\theta_{\lambda}b_{0}))=t_{i}(\sigma(b_{0})\theta_{\lambda})$ for any $i\in I$. For any $x\in\mathbf{f}$ and $i\in I$, it is clear that $\phi_{i}(x)\theta_{\lambda}=\phi_{i}(x\theta_{\lambda})$. By (12), we have ${{}_{i}r}(x)\theta_{\lambda}={{}_{i}r}(x\theta_{\lambda})$. Hence the right multiplication by $\theta_{\lambda}$ commutes with the actions of $\phi_{i},\epsilon_{i}$ for any $i\in I$, and so it commutes with the Kashiwara’s operators $\tilde{\phi_{i}},\tilde{\epsilon}_{i}$ on $\mathbf{f},\mathbf{f}\theta_{\lambda}\mathbf{f}$ for any $i\in I$. By [13, Corollary 17.3.7] and Lemma 5.1, we have $t_{i}(\sigma(b_{0}))=t_{i}(\sigma(b_{0})\theta_{\lambda})$ for any $i\in I$, as desired.
(b) Firstly, we prove that $\theta_{\lambda}b_{0}\in\bigcap_{i\in I}\tilde{\mathbf{B}}_{i,0}$ for any $b_{0}\in\sigma(\mathbf{B}(\lambda))$. Suppose $\theta_{\lambda}b_{0}\in\tilde{\mathbf{B}}_{j,\geqslant 1}$ for some $j\in I$. By [13, Section 14.3.1], we have $t_{j}^{\sigma}(\sigma(\theta_{\lambda}b_{0}))=t_{j}(\theta_{\lambda}b_{0})\geqslant 1$, and so $\sigma(b_{0})\theta_{\lambda}=\sigma(b_{0})\sigma(\theta_{\lambda})=\sigma(\theta_{\lambda}b_{0})\in\tilde{\mathbf{f}}\theta_{j}\subset\sum_{i\in I}\tilde{\mathbf{f}}\theta_{i}$. By Proposition 3.4, we know that $\sigma(b_{0})\in\sum_{i\in I}\mathbf{f}\theta_{i}^{\langle i,\lambda\rangle+1}$, a contradiction to $b_{0}\in\sigma(\mathbf{B}(\lambda))$. Hence $\theta_{\lambda}\cdot$ can be restricted to be an injective map $\sigma(\mathbf{B}(\lambda))\rightarrow\bigcap_{i\in I}\tilde{\mathbf{B}}_{i,0}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$. Secondly, for any $b\in\bigcap_{i\in I}\tilde{\mathbf{B}}_{i,0}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$, we use induction on $\mathrm{tr}\,|b|$ to prove that there exists $b_{0}\in\mathbf{B}$ such that $b=\theta_{\lambda}b_{0}$.
@@
$\bullet$ If $t_{i}^{\sigma}(b)>0$ for some $i\in I$, by Theorem 2.6 and Lemma 4.2, there exists a unique $b^{\prime}\in\tilde{\mathbf{B}}_{i,0}^{\sigma}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ such that $b^{\prime}\theta_{i}^{(t_{i}^{\sigma}(b))}=b+\sum_{b^{\prime\prime}\in\tilde{\mathbf{B}}_{i,\geqslant t_{i}^{\sigma}(b)+1}^{\sigma}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})}c_{b^{\prime\prime}}b^{\prime\prime}$. We claim that $b^{\prime}\in\bigcap_{j\in I}\tilde{\mathbf{B}}_{j,0}$. Indeed, suppose that $t_{j}(b^{\prime})>0$ for some $j\in I$. By Theorem 2.6 and Lemma 4.2, there exists a unique $b^{\prime}_{1}\in\tilde{\mathbf{B}}_{j,0}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ such that $\theta_{j}^{t_{j}(b^{\prime})}b^{\prime}_{1}=b^{\prime}+\sum_{b^{\prime\prime}_{1}\in\tilde{\mathbf{B}}_{j,\geqslant t_{j}(b^{\prime})}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})}c^{\prime}_{b^{\prime\prime}_{1}}b^{\prime\prime}_{1}$. Then we have
$\theta_{j}^{t_{j}(b^{\prime})}b^{\prime}_{1}\theta_{i}^{(t_{i}^{\sigma}(b))}=$ $b^{\prime}\theta_{i}^{(t_{i}^{\sigma}(b))}+\sum_{b^{\prime\prime}_{1}\in\tilde{\mathbf{B}}_{j,\geqslant t_{j}(b^{\prime})}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})}c^{\prime}_{b^{\prime\prime}_{1}}b^{\prime\prime}_{1}\theta_{i}^{(t_{i}^{\sigma}(b))}$
$=$ $b+\sum_{b^{\prime\prime}\in\tilde{\mathbf{B}}_{i,\geqslant t_{i}^{\sigma}(b)+1}^{\sigma}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})}c_{b^{\prime\prime}}b^{\prime\prime}+\sum_{b^{\prime\prime}_{1}\in\tilde{\mathbf{B}}_{j,\geqslant t_{j}(b^{\prime})}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})}c^{\prime}_{b^{\prime\prime}_{1}}b^{\prime\prime}_{1}\theta_{i}^{(t_{i}^{\sigma}(b))}.$
@@
Notice that $c_{b^{\prime\prime}},c^{\prime}_{b^{\prime\prime}_{1}}\in\mathbb{N}[v,v^{-1}]$ and $b^{\prime\prime}_{1}\theta_{i}^{(t_{i}^{\sigma}(b))}$ is a $\mathbb{N}[v,v^{-1}]$-linear combination of $\tilde{\mathbf{B}}$. Hence $b$ appears in $\theta_{j}^{t_{j}(b^{\prime})}b^{\prime}_{1}\theta_{i}^{(t_{i}^{\sigma}(b))}$. By Corollary 2.7, we have $t_{j}(b)\geqslant t_{j}(b^{\prime})>0$, a contradiction. Notice that $\mathrm{tr}\,|b^{\prime}|=\mathrm{tr}\,|b|-t_{i}^{\sigma}(b)$, by the inductive hypothesis, there exists $b^{\prime}_{0}\in\mathbf{B}$ such that $b^{\prime}=\theta_{\lambda}b^{\prime}_{0}$. By (a), we have $t_{i}^{\sigma}(b^{\prime}_{0})=t_{i}^{\sigma}(b^{\prime})=0$. By Theorem 2.6, there exists a unique $b_{0}\in\mathbf{B}_{i,t_{i}^{\sigma}(b)}^{\sigma}$ such that $b^{\prime}_{0}\theta_{i}^{(t_{i}^{\sigma}(b))}=b_{0}+\sum_{b^{\prime\prime}_{0}\in\mathbf{B}_{i,\geqslant t_{i}^{\sigma}(b)+1}^{\sigma}}c^{\prime\prime}_{b^{\prime\prime}_{0}}b^{\prime\prime}_{0}$. Hence
$b-\theta_{\lambda}b_{0}=\sum_{b^{\prime\prime}_{0}\in\mathbf{B}_{i,\geqslant t_{i}^{\sigma}(b)+1}^{\sigma}}c^{\prime\prime}_{b^{\prime\prime}_{0}}\theta_{\lambda}b^{\prime\prime}_{0}-\sum_{b^{\prime\prime}\in\tilde{\mathbf{B}}_{i,\geqslant t_{i}^{\sigma}(b)+1}^{\sigma}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})}c_{b^{\prime\prime}}b^{\prime\prime}.$
@@
By (a), we have $\theta_{\lambda}b_{0}\in\tilde{\mathbf{B}}_{i,t_{i}^{\sigma}(b)}^{\sigma}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$ and $\theta_{\lambda}b^{\prime\prime}_{0}\in\tilde{\mathbf{B}}_{i,\geqslant t_{i}^{\sigma}(b)+1}^{\sigma}\cap\tilde{\mathbf{B}}(\mathbf{f}\theta_{\lambda}\mathbf{f})$. Hence we have $b=\theta_{\lambda}b_{0}$, as desired. It remains to prove that $b_{0}\in\sigma(\mathbf{B}(\lambda))$. Suppose that $t_{j}(b_{0})\geqslant\langle j,\lambda\rangle+1$ for some $j\in I$, since $i^{\prime}\cdot j=-\delta_{i,j}$ for any $i\in I$, by the quantum Serre relations and [13, Lemma 42.1.2(b)], we have